# The Resolution of the Gauge Problem of Cosmology provides Insight into the Formation of the First Structures in the Universe

### v2

@@ A.1 Evolution Equation for Energy Density Perturbations
In a manner analogous to the approach undertaken in Step 1, the explicit appearance of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (129). Differentiating Eq. (129e) with respect to time and subsequently eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{\scriptscriptstyle(1)}$ and $\vartheta_{\scriptscriptstyle(1)}$ with the help of Eqs. (129a)–(129c), results in the following equation:
@@
It can be seen that the conservation laws (10c) and (10d) and the expressions (4a) lead to the following conclusion:
@@
It follows from expressions (137) and (138) that the right-hand side of Eq. (135) does not explicitly contain the gauge-dependent quantities $\varepsilon_{\scriptscriptstyle(1)}$ and $n_{\scriptscriptstyle(1)}$. Consequently, it is gauge-invariant. Therefore, equation (135) governs the evolution of $\varepsilon^{{\textnormal{phys}}}_{{\scriptscriptstyle(1)}}$.
The proof of the equality (137) is straightforward but requires significant computational effort. The algebraic task was performed using the computer algebra system Maxima [44].
The final step in this process will be deriving an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. The linearized equations (123), from which the evolution equations are derived, are of fourth order. From this system, a second-order equation (135) for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ has been derived. It thus follows that the remaining system from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order. As the gauge-invariant quantities, namely, $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been employed, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated. Consequently, only a first-order evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived. In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the expression (138), which contains $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. By differentiating the left-hand side of equation (138) with respect to time and employing the background equations (10c) and (10d), the first-order equations (123a) and (123b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
@@
A.2 Evolution Equations for Density Fluctuations
First Eq. (60a) will be derived. Upon substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ in Eq. (135) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$, the result is
@@ A.2 Evolution Equations for Density Fluctuations
The coefficients $a_{1}$, $a_{2}$ and $a_{3}$ are given by (136). With the aid of the computational software Maxima [44], the coefficients (61) of the main text have been calculated.
@@
By employing the equality (138) and substituting the expression (141) into Eq. (A.1), the following is obtained:
@@
Derivation of Eq. (60b) of the main text was accomplished through the utilization of Eq. (10d).
@@ Appendix B Verifying the Evolution Equations for Scalar Perturbations
The present paper does not provide a derivation of the linearized equations (12). A demonstration of the validity of the system of equations (37), which is derived from the aforementioned linearized equations and is fundamental to the new perturbation theory, will therefore be given via three different methods.
The initial step in this process is to utilize the well-established property of the Einstein equations and conservation laws. This implies that if the constraint equations are satisfied at an initial time, then the solutions of the dynamical equations and conservation laws will satisfy the constraint equations for all subsequent times. A more detailed discussion can be found in Weinberg’s book [31], in Sec. 7.5 on the Cauchy problem. This is also true in reverse: if solutions of the constraint equations and conservation laws can be found for all times, they must necessarily satisfy the dynamical equations. The system (37) comprises the conservation laws and constraint equations, and consists of four first-order ordinary differential equations and one algebraic equation. It thus follows that the system (37), can be solved without the use of the dynamical equations, since the latter are inherently satisfied by the solution of the aforementioned system. Accordingly, the latter of the two observations will be employed to substantiate the validity of equations (37).
The off-diagonal ($i\neq j$) equations of (12c) are not associated with scalar perturbations. For the diagonal ($i=j$) equations, it is sufficient to consider the contraction of the dynamical equations (12c):
@@
Eliminating the quantity $\dot{h}^{k}{}_{k}$ with the help of (19), yields the following result:
@@
Using the energy density constraint equation (37a) to eliminate the second term of (144), results in the following equation:
@@
This equation is identical to the time derivative of the energy density constraint equation (37a), as will now be demonstrated. Differentiation of (37a) with respect to time yields
@@
Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (10a)–(10c), the momentum constraint equation (37b) and the energy density conservation law (37c), respectively, yields the dynamical equation (145). Consequently, Eqs. (37a)–(37c) are correct.
As indicated in the aforementioned verification, the correctness of the equations (37d) and (37e) has not yet been verified. This will now be addressed by employing the gauge modes. It is observed that Eqs. (37) remain invariant under the gauge transformations (1) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (6). Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are known to be gauge-dependent. It is therefore evident that the gauge modes
$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,\theta,p,$ (147a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}$ =~$=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
@@
must be solutions to Eqs. (37). The gauge modes (147a) are derived directly from (3), with the understanding that $\xi^{0}=\psi$. The gauge mode (147b) is the covariant divergence of (28). The gauge mode (147c) is derived by substituting the gauge modes (147a)–(147b) into the perturbed Friedmann equation (37a):
@@
Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (10a), and Eq. (10b), one arrives at the gauge mode (147c). It will now be demonstrated that the gauge modes given by (147) are indeed solutions of Eqs. (37b)–(37e). In order to execute this task, the first step entails deriving the time derivatives of the gauge modes:
@@
where Eqs. (10b)–(10d) and Eq. (44) have been used. Subsequent to the aforementioned preparation, the verification of the accuracy of the equations (37b)–(37e) can be undertaken. First, by substituting the gauge modes, namely $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq. (37b), the result is an equation from which the variable $\psi$ disappears. Next, dividing the resulting equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying by $\tfrac{3}{2}H$, one arrives at the time derivative of Eq. (10a). Second, the substitution of the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37c) results in the elimination of the term with $\dot{H}$. Next, dividing, the resulting equation by $\psi$ and employing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by utilizing Eq. (10c) yields an identity. Third, by substituting the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37d), and utilizing the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq. (10c) one arrives at an identity. Finally, upon substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq. (37e) the term with $\dot{H}$ vanishes. Using Eq. (10d), one arrives at an identity. It has been demonstrated that the gauge modes (147) satisfy Eqs. (37b)–(37e). Consequently, it can be concluded that these equations are accurate. Given the established validity of Eq. (37a), it can be deduced that the system (37) is also accurate.
As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (152) for the gauge-dependent density fluctuation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs. (37). Of particular note is the fact that the well-known evolution equations for $\delta$ are recovered in the cases of a radiation-dominated universe (154) and a universe after decoupling of matter and radiation (156).
@@
Appendix C Conventional Evolution Equation for Cosmological Density Fluctuations
The usual evolution equation for density perturbations in a flat, ${{}^{3}\!}R_{{\scriptscriptstyle(0)}}=0$, flrw universe is given by
$\ddot{\delta}+2H\dot{\delta}-\left(\beta^{2}\frac{\nabla^{2}}{a^{2}}+\tfrac{1}{2}\kappa\varepsilon_{\scriptscriptstyle(0)}(1+w)(1+3w)\right)\delta=0,$ (150)
where $w:=p_{{\scriptscriptstyle(0)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ and $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$. The derivation of this equation has four primary objectives. Firstly, it demonstrates the correctness of the system (37), at least in the context of a flat flrw universe. Secondly, it is demonstrated that the standard equation (150) is incomplete. That is, its solution is a linear combination of only one physical mode and one gauge mode, independent of the scale of a fluctuation. Consequently, Eq. (150) does not give an adequate description of cosmological density perturbations. Thirdly, it is shown that the novel perturbation theory (60) provides all well-known physical solutions for large-scale perturbations in a flat flrw universe. Specifically, it produces the sets $\{t,t^{1/2}\}$ for the radiation-dominated universe and $\{t^{2/3},t^{-5/3}\}$ for the universe after decoupling of matter and radiation. Finally, it will be demonstrated that Eq. (150), when derived from the Newtonian theory of gravitation, yields solutions that remain dependent on the selection of the coordinates.
In order to derive equation (150), it is necessary to utilize the background equations (71) and the equations (123) for scalar perturbations. These two systems of equations are complete, since they contain all conservation laws and constraint equations. In this appendix, it is assumed that $c=1$.
Equation (150) has been derived from an equation of state $p=p(\varepsilon)$ in the existing literature. Therefore, the particle number conservation law (71c) and its linearized counterpart (123b) are not required. Substituting $p_{n}=0$ into (13) and (40) yields $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$. According to the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs. (123) can be rewritten as follows:

### v3

@@ A.1 Evolution Equation for Energy Density Perturbations
In a manner analogous to the approach undertaken in Step 1, the explicit occurrence of $\vartheta_{{\scriptscriptstyle(1)}}$ is removed from the system of equations (128). Differentiating Eq. (128e) with respect to time and subsequently eliminating the time derivatives of $\varepsilon_{{\scriptscriptstyle(1)}}$, $n_{\scriptscriptstyle(1)}$ and $\vartheta_{\scriptscriptstyle(1)}$ with the help of Eqs. (128a)–(128c), results in the following equation:
@@
The proof of this equality is straightforward but requires significant computational effort. The algebraic task was performed using the computer algebra system Maxima [52]. The conservation laws (11c) and (11d) and the expressions (6a) yield:
@@
Substituting (135) into (134) and using (136), we arrive at the evolution equation for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$:
[…]
where the coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by
The derivation of Eq. (137) shows that both the perturbation ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ (19) of the spatial curvature and the divergence $\vartheta_{{\scriptscriptstyle(1)}}$ (20) of the fluid three-vector are included in this equation.
The final step in this process will be deriving an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$. The linearized equations (122), from which the evolution equations are derived, are of fourth order. From this system, a second-order equation (137) for $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ has been derived. It thus follows that the remaining system (133a)–(133b), and (133e) from which an evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived is at most of second order. As the gauge-invariant quantities, namely $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ and $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, have been used, one degree of freedom, specifically the gauge function $\psi(\bm{x})$ in (147), has been eliminated. Consequently, only a first-order evolution equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$ can be derived. In place of deriving an equation for $n^{\textnormal{phys}}_{\scriptscriptstyle(1)}$, an evolution equation will be derived for the right-hand side of expression (136). By differentiating the left-hand side of this expression with respect to time and using the background equations (11c) and (11d), the first-order equations (122a) and (122b) and the definitions $w:=p_{\scriptscriptstyle(0)}/\varepsilon_{\scriptscriptstyle(0)}$ and $\beta^{2}:=\dot{p}_{\scriptscriptstyle(0)}/\dot{\varepsilon}_{\scriptscriptstyle(0)}$, the following result is obtained:
@@
A.2 Perturbation Theory
First Eq. (62a) will be derived. Substituting the expression $\varepsilon^{\textnormal{phys}}_{\scriptscriptstyle(1)}=\varepsilon_{\scriptscriptstyle(0)}\delta_{\varepsilon}$ into Eq. (137) and dividing by $\varepsilon_{\scriptscriptstyle(0)}$ yields the following result:
@@ A.2 Perturbation Theory
The coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by (138). With the aid of the computational software Maxima [52], the coefficients (63) of the main text have been calculated.
@@
By using the equality (136) and substituting the expression (141) into Eq. (A.1), the following is obtained:
@@
Using Eq. (11d) yields Eq. (62b) of the main text.
@@ Appendix B Verifying the Evolution Equations for Scalar Perturbations
This article does not present a derivation of the linearized equations (13), from which the system of equations (37) for scalar perturbations is derived. Since the latter system is key to the cosmological perturbation theory developed in this article, we will demonstrate its validity. This will be accomplished using three different methods.
The first step in this process is to apply a well-known property of Einstein’s equations and conservation laws. This property stems from the fact that the constraint equations contain only first-order time derivatives of the metric tensor, while the dynamical equations contain second-order time derivatives. It implies that if the constraint equations are satisfied at an initial time, then the solutions to the dynamical equations and conservation laws will also satisfy the constraint equations at all subsequent times. A detailed discussion can be found in Sec. 7.5 of Weinberg’s textbook [42] on the Cauchy problem. The reverse is also true: if solutions to the constraint equations and conservation laws can be found for all times, then they necessarily satisfy the dynamical equations. We will now use this latter property to validate the system of equations (37) for scalar perturbations.
The system (37) encompasses the conservation laws and constraint equations. It consists of four first-order ordinary differential equations and one algebraic equation. Therefore, it can be solved, and its solutions automatically satisfy the system of dynamical equations. This fact will be used to prove the correctness of Eqs. (37).
For $i\neq j$, the dynamical equations (13c) are not coupled to the scalar perturbations. For the case of $i=j$, it is sufficient to consider the contraction of (13c):
@@
Using (20), yields
@@
Eliminating the second term of equation (144) using the energy density constraint equation (37a) results in
@@
As will now be shown, this equation is identical to the time derivative of the energy density constraint equation (37a). Differentiating equation (37a) with respect to time yields
@@
Eliminating the time derivatives $\dot{H}$, ${{}^{3}\!}\dot{R}_{\scriptscriptstyle(1)}$, and $\dot{\varepsilon}_{\scriptscriptstyle(1)}$ with the help of (11a)–(11c), the momentum constraint equation (37b) and the energy density conservation law (37c) yields the dynamical equation (145). Consequently, Eqs. (37a)–(37c) are correct.
We will now continue with the second method of verifying Eqs. (37). Equations (37d) and (37e) have not yet been verified by the above procedure. This will now be addressed by using the gauge modes. Equations (37) are invariant under the gauge transformations (3) with the gauge functions $\xi^{\mu}(t,\bm{x})$ given by (9). Furthermore, the quantities $\varepsilon_{\scriptscriptstyle(1)}$, $n_{\scriptscriptstyle(1)}$, $\theta_{\scriptscriptstyle(1)}$, $\vartheta_{\scriptscriptstyle(1)}$, and ${{}^{3}\!}R_{\scriptscriptstyle(1)}$ are gauge-dependent. It is therefore evident that the gauge modes
$\hat{S}_{\scriptscriptstyle(1)}$ $=\psi\dot{S}_{\scriptscriptstyle(0)},\quad S=\varepsilon,n,T,p,\theta,$ (147a)
$\hat{\vartheta}_{\scriptscriptstyle(1)}$ $=-\frac{\tilde{\nabla}^{2}\psi}{a^{2}},$ (147b)
@@
must be solutions to Eqs. (37). The gauge modes (147a) follow from (5), with the understanding that $\xi^{0}=\psi(\bm{x})$. The gauge mode (147b) is the covariant divergence of (29). As shown above, the linearized Friedmann equation (37a) is correct. Therefore, it can be used to derive the gauge mode (147c). Substituting the gauge modes (147a)–(147b) into the linearized Friedmann equation results in
@@
Using that $\theta_{{\scriptscriptstyle(0)}}=3H$, the Friedmann equation (11a), Eq. (11b), and the energy density conservation law (11c) yields the gauge mode (147c). Expressions (147) imply that the quantities $S_{{\scriptscriptstyle(1)}}$, $\vartheta_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}R_{{\scriptscriptstyle(1)}}$ are invariant under spatial gauge transformations, i.e., transformations for which the condition $\psi(\bm{x})=0$ holds. However, they are not invariant under space-time gauge transformations.
It will now be demonstrated that the gauge modes given by (147) are solutions to Eqs. (37b)–(37e). The first step to doing so is deriving the time derivatives of the gauge modes:
@@
where Eqs. (11b)–(11d) and Eq. (44) have been used. The validation of the Eqs.  (37b)–(37e) can now be conducted.
First, substitute the gauge modes $\hat{\theta}_{{\scriptscriptstyle(1)}}$, $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and ${{}^{3}\!}\hat{R}_{{\scriptscriptstyle(1)}}$ into Eq. (37b), to obtain an equation from which the gauge function $\psi$ disappears. Next, dividing this equation by $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$ and multiplying the result by $\tfrac{3}{2}H$ yields the time derivative of Eq. (11a). Second, substitute the gauge modes $\hat{\varepsilon}_{{\scriptscriptstyle(1)}}$, $\hat{\theta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37c) eliminates the term with $\dot{H}$. Next, dividing the resulting equation by $\psi$ and using the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ followed by the use of Eq. (11c) yields an identity. Third, substitute the gauge modes $\hat{\vartheta}_{{\scriptscriptstyle(1)}}$, and $\hat{p}_{{\scriptscriptstyle(1)}}$ into Eq. (37d), and using the definition $\beta^{2}:=\dot{p}_{{\scriptscriptstyle(0)}}/\dot{\varepsilon}_{{\scriptscriptstyle(0)}}$ and subsequently using Eq. (11c) one arrives at an identity. Finally, by substituting the gauge modes $\hat{n}_{{\scriptscriptstyle(1)}}$ and $\hat{\theta}_{{\scriptscriptstyle(1)}}$ into Eq. (37e) the term with $\dot{H}$ vanishes. Using Eq. (11d), one arrives at an identity. It has been demonstrated that the gauge modes (147) satisfy Eqs. (37b)–(37e). Consequently, it can be concluded that these equations are correct. Given the established validity of Eq. (37a), it can be deduced that the system (37) is accurate as well.
As a final check, it is observed in Appendix C that for a flat flrw universe, the evolution equations (151) for the gauge-dependent relative perturbation $\delta:=\varepsilon_{{\scriptscriptstyle(1)}}/\varepsilon_{{\scriptscriptstyle(0)}}$ are obtained from Eqs. (37). In fact, the well-known evolution equations for $\delta$ are obtained in the cases of a radiation-dominated universe (154) and a universe after matter and radiation decouple (157).
@@
Appendix C Conventional Evolution Equation for Relative Density Perturbations
In this appendix the usual perturbation equations for relative perturbations in a flat, ${{}^{3}\!}R_{\scriptscriptstyle(0)}=0$, flrw universe is compared with the perturbation theory developed in this article. To that end, the evolution equation for the relative perturbation $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$ is derived using the systems of background equations (75) and evolution equations (122) for scalar perturbations. In this appendix, we assume $c=1$ and $\Lambda=0$.
In the existing literature, the particle number density is usually not taken into account. Therefore, we take the equation of state for the pressure $p=p(\varepsilon)$, so that $p_{n}=0$}. From (14) and (40) we find that $p_{{\scriptscriptstyle(1)}}=p_{\varepsilon}\varepsilon_{{\scriptscriptstyle(1)}}=\beta^{2}\varepsilon_{{\scriptscriptstyle(1)}}$. Using the definition $\delta:=\varepsilon_{\scriptscriptstyle(1)}/\varepsilon_{\scriptscriptstyle(0)}$, Eqs. (122) can be rewritten as follows: